# SpeechPrune: Context-aware Token Pruning for Speech Information Retrieval

### v1

@@ Abstract
We introduce Speech Information Retrieval (SIR), a new long-context task for Speech Large Language Models (Speech LLMs), and present SPIRAL, a 1,012-sample benchmark testing models’ ability to extract critical details from approximately 90-second spoken inputs. While current Speech LLMs excel at short-form tasks, they struggle with the computational and representational demands of longer audio sequences. To address this limitation, we propose SpeechPrune, a training-free token pruning strategy that uses speech-text similarity and approximated attention scores to efficiently discard irrelevant tokens. In SPIRAL, SpeechPrune achieves accuracy improvements of 29% and up to 47% over the original model and the random pruning model at a pruning rate of 20%, respectively. SpeechPrune can maintain network performance even at a pruning level of 80%. This approach highlights the potential of token-level pruning for efficient and scalable long-form speech understanding.
@@ I Introduction
Speech Large Language Models (Speech LLMs) represent a significant advancement in speech language understanding and processing, as they leverage contextual reasoning capabilities of large language models to process audio inputs [bu2024roadmap]. Unlike traditional cascaded pipelines, where automatic speech recognition (ASR) and language modeling are handled by separate modules, Speech LLMs unify audio processing, cross-modal fusion, and language modeling in a single architecture [peng2024survey]. These unified models can perform multiple tasks like speech recognition, speech translation, speaker identification and emotion recognition, while maintaining end-to-end trainability [chu2024qwen2, DiVA, zhang2023speechgpt, zhan2024anygpt].
Despite the broad applications of Speech LLMs, there is one desirable functionality for these models that remains largely unexplored by existing work. Specifically, it is the capability of extracting crucial information within long-context audio, which we term Speech Information Retrieval (SIR). SIR is particularly relevant to real-world scenarios, which often require extracting key information from extended audio content, such as meetings, lectures, interviews, and customer service calls. For instance, the user may want the model (as an AI assistant) to accurately note down the time for a future event mentioned in a long conversation, so as to help them optimize their schedule. While straightforward to be accomplished by us humans, SIR is non-trivial and challenging for Speech LLMs. First, the target information is very likely to exist only in one short audio segment among the full, extensively long audio inputs. Precisely recognizing the relevant parts and ignoring the irrelevant parts is intuitively challenging for the models. Second, as we will discuss later, a more prohibitive limitation for Speech LLMs to perform SIR is their significant computational inefficiency when processing long audio token sequences.
@@
Figure 1: The proposed SpeechPrune, with two levels of token pruning.
To address the limitation, our second technical contribution is SpeechPrune, a training-free token pruning method that enables off-the-shelf Speech LLMs to handle lengthy audio input efficiently and effectively. SpeechPrune features a two-stage process, where it first gets rid of semantically irrelevant speech tokens by examining the cosine similarity between speech and text token embeddings, and then further selects the most important tokens by approximating token importance with binarized attention weights from the first layer. This plug-and-play approach maintains semantic fidelity while substantially reducing computational overhead, making the processing of long audio inputs possible without any additional training upon pre-trained models. Our SpeechPrune, which to the best of our knowledge is the first token pruning method for Speech LLMs, achieves nearly 29% (and 47%) higher accuracy than the original model (and the random pruning baseline) at a 20% pruning rate and sustains performance even when pruning 80% of the input on our SPIRAL benchmark.
@@ II-B Benchmark Construction
We introduce SPIRAL (Speech Information Retrieval And Lookup), a novel benchmark designed to evaluate Speech LLMs’ ability to process long and realistic spoken inputs.
The samples in our dataset features three representative scenarios, including lectures, meetings, and daily conversations. Within each scenario, there are various fine-grained and specific topics that ultimately form a diverse and hierarchical topic structure for SPIRAL (see Fig. 4-6 in Appendix for details). Unlike existing approaches that simply apply speech synthesis to transform text datasets into speech datasets, we specifically design our data to reflect the unique characteristics of oral communication by considering two steps, namely transcript generation and speech sample synthesis, in our construction.
@@
    Topic Curation: A comprehensive array of topics is meticulously selected to capture the breadth and complexity of human interactions. These topics are detailed in Appendix -A1.
@@
    Dialogue Generation: Using GPT-4o, we generate multi-turn dialogues incorporating oral speech elements (fillers like “uh” and “oh”) to enhance authenticity. The detailed prompt is provided in Appendix -A2. Multiple-choice questions are generated for evaluation purposes.
@@
    Speech Generation: Using StyleTTS 2, we generate speech with fixed diffusion steps and embedding scale parameters. Dialogue turns are concatenated to create continuous speech while preserving conversational flow. The detailed parmaeters used are listed in Appendix B.
The complete SPIRAL dataset is open-source [^1] https://github.com/linyueqian/SPIRAL_Dataset, facilitating further research in SIR tasks. In addition, we propose SPIRAL-H, a challenging subset containing 401 cases where the original Qwen-2 Audio model in the experiment completely fails with 0% accuracy. We show some examples in Appendix -A3.
@@ II-C Quality Assessment
The generated SPIRAL dataset contain 1,012 samples, with an average duration of 87.89 seconds.
To assess the quality of our generated SPIRAL dataset, we evaluate the synthesized samples using two complementary metrics: automatic speech recognition accuracy via Whisper-v3-large [whisper], which achieves a word error rate of 0.0389, and perceptual quality via UTMOS-22 [saeki22c_interspeech], a widely used surrogate objective metric of mean opinion score (MOS), yielding a predicted MOS of 3.91 in a five-point-scale. These metrics respectively quantify the transcription accuracy of the speech content by a state-of-the-art recognition system and the naturalness/human-likeness of the speech, as evaluated by a perceptual quality model. SPIRAL demonstrates strong performance in both metrics.
@@ III-A Preliminaries
Audio Encoder Speech LLMs typically consist of an audio encoder (such as Whisper [whisper]) which transforms raw audio with high sampling rates into lower-dimensional embeddings. Taking Whisper as an example, an audio input (with maximum length) is first processed and transformed into a 80-channel melspectrogram in the time-frequency domain. This 80-channel melspectrogram, generated with a window size of 25 ms and a hop size of 10 ms, is then fed into the Transformer-based encoder. A pooling layer with a stride of two follows to reduce the length of the audio representation. As a result, each frame of the encoder output approximately corresponds to a 40ms segment of the original audio signal. Thus, a 30-second audio yields 750 encoding embeddings. This temporal correspondence between audio frames and encoder outputs provides a natural foundation for our frame-level pruning strategy, as we can leverage the inherent structure of how speech information is encoded to maintain temporal coherence during pruning.
@@ III-B SpeechPrune Methodology
We propose a two-stage token pruning approach, as shown in Fig. 1 and the following parts.
First Level Pruning by Token-Text Similarity The first stage utilizes the correlation between audio and text tokens to identify semantically important audio segments. Recent research has shown that such audio-text token alignment enables effective cross-modal reasoning in speech language models [DiVA]. More formally, we process the input to get speech embedding $\mathbf{S}\in\mathbb{R}^{N\times D}$ and text embedding $\mathbf{T}\in\mathbb{R}^{L\times D}$, where $N$ is the number of speech tokens before pruning, $L$ is the number of text tokens, and $D$ is the embedding dimensionality. Here we only consider real text query as $T$ and exclude system prompt and special tokens. The token-level similarity matrix $\mathbf{F}\in\mathbb{R}^{N\times L}$ between speech and text tokens is computed using cosine similarity:
@@
The speech token remained after first level pruning is:
@@
Second Level Pruning by Binarized Attention Estimation Building on the first-level pruning results, we introduce a second pruning stage to further select important tokens based on approximated attention scores. This stage exclusively focuses on speech tokens, as the text-speech relationships have already been captured in the first pruning stage, enabling efficient modeling of internal dependencies within speech segments while minimizing computational overhead. The second stage utilizes the binarized attention from the network’s first transformer layer. Specifically, we compute the scores using the signed binarized Query and Key weights, and also the pruned speech embeddings:
@@
$\mathbf{A}=\operatorname{softmax}(\frac{\mathbf{Q}^{\prime}{\mathbf{K}^{\prime}}^{\top}}{\sqrt{d_{k}}})\in\mathbb{R}^{n\times n}.$ (12)
@@ IV Experiments
We conduct our main experiments using Qwen-2 Audio [chu2024qwen2], a state-of-the-art Speech LLM with extensive speech understanding task coverage. Our primary results are presented in Section IV-A, with qualitative analyses discussed in Section IV-B. Additionally, we perform ablation studies examining the performance impact of each pruning level in Section IV-C. Finally, Section IV-D demonstrates the generalizability of our proposed method across different models and benchmarks.
@@ IV-A Main Experiments
Original - $12.2$ $779$ $13.40$ $0.19$ $60.38$% $0\text{\%}$
RAP $0.2$ $10.06$ $662$ $13.32$ $0.15$ $42.49\text{\%}$ $21.45\text{\%}$
RAC $65.71\text{\%}$ $48.13\text{\%}$
Ours $\mathbf{89.23\textbf{\%}}$ $\mathbf{81.64\textbf{\%}}$
RAP $0.4$ $7.93$ $511$ $13.24$ $0.11$ $42.89\text{\%}$ $22.19\text{\%}$
RAC $62.45\text{\%}$ $41.90\text{\%}$
Ours $\mathbf{85.97\text{\%}}$ $\mathbf{76.43\text{\%}}$
RAP $0.6$ $5.79$ $419$ $13.17$ 0.07 $42.39\text{\%}$ $21.45\text{\%}$
RAC $58.20\text{\%}$ $35.41\text{\%}$
Ours $\mathbf{75.89\text{\%}}$ $\mathbf{63.77\text{\%}}$
RAP $0.8$ $3.66$ $278$ $13.09$ $0.04$ $45.26\text{\%}$ $23.19\text{\%}$
RAC $55.83\text{\%}$ $33.67\text{\%}$
Ours $\mathbf{62.45\text{\%}}$ $\mathbf{46.15\text{\%}}$
Setup We evaluate our method using Qwen-2 Audio, comparing our SpeechPrune method against several baselines, comparing our two-level pruning strategy (SpeechPrune) against three baselines: (1) Original: full audio trimmed at 30 seconds (750 tokens); (2) RAP: random audio pruning that selects non-contiguous segments to reach target rate; and (3) RAC: random audio cropping that selects a single contiguous segment at target rate. Our SpeechPrune’s two-level pruning strategy is set as follows: the first level prunes the input tokens to match the original method’s input length (which is 750 tokens), while the second level removes additional tokens according to the specified pruning rate. We assess computational efficiency via TFLOPS [^2] Calculate by calflops: https://github.com/MrYxJ/calculate-flops.pytorch, prefill time measured on a Quadro RTX6000 GPU, and memory usage (total and activations), estimated with LLM-Viewer [yuan2024llm].
@@ IV-C Ablation Studies
To evaluate the effectiveness of our two-level pruning approach, we conduct ablation studies on the SPIRAL-H dataset. We examine three variants of our method: using only the first level pruning, using only the second level pruning, and the complete two-stage approach. Fig. 3 presents the performance comparison across different pruning rates. When using the complete set of unpruned input tokens, the model achieves an accuracy of 43.6%. The combined approach consistently outperforms both individual pruning levels across most pruning rates, achieving peak performance of $81.64\%$ at $0.2$ pruning rate compared to $48.13\%$ and $72.45\%$ for first level and second level only, respectively. This significant improvement over the original model’s $0\%$ accuracy on SPIRAL-H indicates that our pruning strategy not only reduces computational cost but also enhances the model’s ability to identify and process critical information. Second, we observe interesting behavioral patterns for each variant: the first level only approach shows relatively stable but lower performance ($45$-$55\%$), while the second level only method starts with higher accuracy but degrades more rapidly as pruning rate increases. Finally, the combined approach exhibits the most robust performance, maintaining superior accuracy until around $0.7$ pruning rate, after which all methods converge to similar performance levels. This suggests that our two-level design effectively leverages complementary information from both token-level similarity and attention patterns, resulting in more robust and efficient pruning even on challenging cases where the original model fails.
Figure 3: Ablation study comparing different pruning strategies on SPIRAL-H dataset. The plot shows the accuracy of three approaches: first level only, second level only, and combined steps. The dotted line shows the accuracy (43.6%) when using the complete, unpruned set of input tokens.
@@ IV-D Generalization Analysis
To evaluate the generalization capability of our method, we test SpeechPrune on both different benchmarks and a different Speech LLM model. For additional benchmarks, we select two representative long-form speech understanding datasets: DREAM-TTS and CN-College-Listen. DREAM-TTS is derived from the text-based dialogue comprehension dataset DREAM [DREAM], converted to speech using state-of-the-art TTS technology with 60 different speakers while maintaining gender consistency by [wang2024audiobench]. CN-College-Listen is sourced from WavLLM’s [hu2024wavllm] test set, comprising English listening comprehension questions from China’s national college entrance examinations. For both datasets, we specifically use test samples that exceed 60 seconds in length to evaluate long-form speech understanding capabilities.

### v2

@@ Abstract
While current Speech Large Language Models (Speech LLMs) excel at short-form tasks, they struggle with the computational and representational demands of longer audio clips. To advance the model’s capabilities with long-form speech, we introduce Speech Information Retrieval (SIR), a long-context task for Speech LLMs, and present SPIRAL, a 1,012-sample benchmark testing models’ ability to extract critical details from long spoken inputs. To overcome the challenges of processing long speech sequences, we propose SpeechPrune, a training-free token pruning strategy that uses speech-text similarity and approximated attention scores to efficiently discard irrelevant tokens. In SPIRAL, SpeechPrune achieves accuracy improvements of 29% and up to 47% over the original model and the random pruning model at a pruning rate of 20%, respectively. SpeechPrune can maintain network performance even at a pruning level of 80%. This highlights the potential of token-level pruning for efficient and scalable long-form speech understanding.
@@ I Introduction
Speech Large Language Models (Speech LLMs) represent a significant advancement in speech-language understanding and processing, as they leverage contextual reasoning capabilities of large language models to process audio inputs. Unlike traditional cascaded pipelines, where automatic speech recognition (ASR) and language modeling are handled by separate modules, Speech LLMs unify audio processing, cross-modal fusion, and language modeling in a single architecture [peng2024survey]. These unified models can perform multiple tasks like speech recognition, speech translation, speaker identification and emotion recognition, while maintaining end-to-end trainability [chu2024qwen2, DiVA, zhang2023speechgpt, zhan2024anygpt].
Despite the broad applications of Speech LLMs, one desirable functionality for these models remains unexplored in existing work. Specifically, it is the capability of extracting crucial information within long-context audio, which we term Speech Information Retrieval (SIR). SIR is particularly relevant to real-world scenarios, which often require extracting key information from extended audio content, such as meetings, lectures, interviews, and customer service calls. For instance, the user may want the model (as an AI assistant) to accurately note down the time for a future event mentioned in a long conversation, so as to help them optimize their schedule. While straightforward to be accomplished by us humans, SIR is non-trivial and challenging for Speech LLMs. First, the target information will likely exist only in one short audio segment among the whole, extensively long audio inputs. Precisely recognizing the relevant parts and ignoring the irrelevant parts is intuitively challenging for the models. Second, as we will discuss later, a more prohibitive limitation for Speech LLMs to perform SIR is their significant computational inefficiency when processing long audio token sequences.
@@
Figure 1: The proposed SpeechPrune, with two phases of token pruning.
To address the limitation, our second technical contribution is SpeechPrune, a training-free token pruning method that enables off-the-shelf Speech LLMs to handle lengthy audio input efficiently and effectively. Unlike existing vision-centric pruning methods (e.g., PruMerge [shang2024llava]) that are incompatible with speech encoders, SpeechPruneis specifically designed to preserve the temporal nature of audio signals. SpeechPrune features a two-phase process, where it first removes semantically irrelevant speech tokens by examining the cosine similarity between speech and text token embeddings, and then further selects the most important tokens by approximating token importance with binarized attention weights from the first layer. This plug-and-play approach maintains semantic fidelity while substantially reducing computational overhead, making the processing of long audio inputs possible without any additional training upon pre-trained models. Our SpeechPrune, which, to the best of our knowledge is the first token pruning method for Speech LLMs, achieves nearly 29% (and 47%) higher accuracy than the original model (and the random pruning baseline) at a 20% pruning rate and sustains performance even when pruning 80% of the input on our SPIRAL benchmark.
@@ II-B Benchmark Construction
We introduce SPIRAL (Speech Information Retrieval And Lookup), a novel benchmark designed to evaluate Speech LLMs’ ability to process long and realistic spoken inputs. The samples in our dataset feature three representative scenarios, including lectures, meetings, and daily conversations. Within each scenario, there are various fine-grained and specific topics that ultimately form a diverse and hierarchical topic structure for SPIRAL. Unlike existing approaches that simply apply speech synthesis to transform text datasets into speech datasets, we specifically design our data to reflect the unique characteristics of oral communication through a systematic two-stage pipeline, namely transcript generation and speech sample synthesis, in our construction.
@@
    Topic Curation: A comprehensive array of topics is meticulously selected to capture the breadth and complexity of human interactions, with hierarchial orgnization to ensure diverse coverage across domains.
@@
    Dialogue Generation: Using GPT-4o, we generate multi-turn dialogues incorporating natural speech elements (fillers like “uh” and “oh”) to enhance authenticity. Our prompt engineering specifically guides the model to create realistic conversational dynamics with variable turn lengths and contextual continuity. Multiple-choice questions are generated for evaluation purposes.
@@
    Speech Generation: Using StyleTTS 2, we generate speech with fixed diffusion steps and embedding scale parameters. Dialogue turns are concatenated to create continuous speech while preserving conversational flow.
The SPIRAL dataset is open-source, [^1] The dataset and its construction details can be accessed through our website. facilitating further research on SIR tasks. In addition, we propose SPIRAL-H, a challenging subset consisting of 401 cases in which the original Qwen-2 Audio model used in our experiments fails completely, achieving 0% accuracy.
@@ II-C Quality Assessment
The generated SPIRAL dataset contains 1,012 samples, with an average duration of 87.89 seconds. To assess the quality of our generated SPIRAL dataset, we evaluate the synthesized samples using two complementary metrics: automatic speech recognition accuracy via Whisper-v3-large [whisper], which achieves a word error rate of 0.0389, and perceptual quality via UTMOS-22 [saeki22c_interspeech], a widely used surrogate objective metric of mean opinion score (MOS), yielding a predicted MOS of 3.91 in a five-point-scale. These metrics respectively quantify the transcription accuracy of the speech content by a state-of-the-art recognition system and the naturalness/human-likeness of the speech, as evaluated by a perceptual quality model. SPIRAL demonstrates strong performance in both metrics.
@@ III-A Preliminaries
Audio Encoder Speech LLMs typically consist of an audio encoder (such as Whisper [whisper]) which transforms raw audio with high sampling rates into lower-dimensional embeddings. Taking Whisper as an example, an audio input (with maximum length) is first processed and transformed into an 80-channel melspectrogram in the time-frequency domain. This 80-channel melspectrogram, generated with a window size of 25 ms and a hop size of 10 ms, is then fed into the Transformer-based encoder. A pooling layer with a stride of two follows to reduce the length of the audio representation. As a result, each frame of the encoder output approximately corresponds to a 40ms segment of the original audio signal. Thus, a 30-second audio yields 750 encoding embeddings. This temporal correspondence between audio frames and encoder outputs provides a natural foundation for our frame-level pruning strategy, as we can leverage the inherent structure of how speech information is encoded to maintain temporal coherence during pruning.
@@ III-B SpeechPrune Methodology
We propose a two-phase token pruning approach, as shown in Fig. 1 and the following parts.
First Phase Pruning by Token-Text Similarity The first phase utilizes the correlation between audio and text tokens to identify semantically important audio segments. Recent research has shown that such audio-text token alignment enables effective cross-modal reasoning in speech-language models [DiVA]. More formally, we process the input to get speech embedding $\mathbf{S}\in\mathbb{R}^{N\times D}$ and text embedding $\mathbf{T}\in\mathbb{R}^{L\times D}$, where $N$ is the number of speech tokens before pruning, $L$ is the number of text tokens, and $D$ is the embedding dimensionality. Here, we only consider real text query as $\mathbf{T}$ and exclude system prompt and special tokens. The token-level similarity matrix $\mathbf{F}\in\mathbb{R}^{N\times L}$ between speech and text tokens is computed using cosine similarity:
@@
The speech token remaining after first phase pruning is:
@@
Second Phase Pruning by Binarized Attention Estimation Building on the first-phase pruning results, we introduce a second pruning phase to further select important tokens based on approximated attention scores. This phase exclusively focuses on speech tokens, as the text-speech relationships have already been captured in the first pruning phase, enabling efficient modeling of internal dependencies within speech segments while minimizing computational overhead. The second phase utilizes the binarized attention from the network’s first transformer layer. Specifically, we compute the scores using the signed binarized Query and Key weights, and also the pruned speech embeddings:
@@
$\mathbf{A}=\operatorname{softmax}(\frac{\mathbf{Q}^{\prime}{\mathbf{K}^{\prime}}^{\top}}{\sqrt{d_{k}}}).$ (12)
@@ IV Experiments
We conduct our main experiments using Qwen-2 Audio [chu2024qwen2], a state-of-the-art Speech LLM with extensive speech understanding task coverage. Our primary results are presented in Section IV-A, with qualitative analyses discussed in Section IV-B. Additionally, we perform ablation studies examining the performance impact of each pruning phase in Section IV-C. Finally, Section IV-D demonstrates the generalizability of our proposed method across different models and benchmarks.
@@ IV-A Main Experiments
Original 12.2 779 13.40 0.19 60.38% 0%
RAP 0.2 10.06 662 13.32 0.15 42.49% 21.45%
RAC 65.71% 48.13%
Ours 89.23% 81.64%
RAP 0.4 7.93 511 13.24 0.11 42.89% 22.19%
RAC 62.45% 41.90%
Ours 85.97% 76.43%
RAP 0.6 5.79 419 13.17 0.07 42.39% 21.45%
RAC 58.20% 35.41%
Ours 75.89% 63.77%
RAP 0.8 3.66 278 13.09 0.04 45.26% 23.19%
RAC 55.83% 33.67%
Ours 62.45% 46.15%
Setup We evaluate our method using Qwen-2 Audio, comparing our SpeechPrune method against several baselines, comparing our two-phase pruning strategy (SpeechPrune) against three baselines: (1) Original: full audio trimmed at 30 seconds (750 tokens); (2) RAP: random audio pruning that selects non-contiguous segments to reach target rate; and (3) RAC: random audio cropping that selects a single contiguous segment at target rate. Our SpeechPrune’s two-phase pruning strategy is set as follows: the first phase prunes the input tokens to match the original method’s input length (which is 750 tokens), while the second phase removes additional tokens according to the specified pruning rate. We evaluate computational efficiency using TFLOPS [^2] Calculated using calflops: https://github.com/MrYxJ/calculate-flops.pytorch, measure prefill time on a Quadro RTX6000 GPU, and assess memory usage (total and activations) using LLM-Viewer [yuan2024llm].
@@ IV-C Ablation Studies
To evaluate the effectiveness of our two-phase pruning approach, we conduct ablation studies on the SPIRAL-H dataset. We examine three variants of our method: using only the first phase pruning, using only the second phase pruning, and the complete two-phase approach. Fig. 3 presents the performance comparison across different pruning rates. When using the complete set of unpruned input tokens, the model achieves an accuracy of 43.6%. The combined approach consistently outperforms both individual pruning phases across most pruning rates, achieving peak performance of $81.64\%$ at $0.2$ pruning rate compared to $48.13\%$ and $72.45\%$ for first phase and second phase only, respectively. This significant improvement over the original model’s $0\%$ accuracy on SPIRAL-H indicates that our pruning strategy not only reduces computational cost but also enhances the model’s ability to identify and process critical information. Second, we observe interesting behavioral patterns for each variant: the first phase only approach shows relatively stable but lower performance ($45$-$55\%$), while the second phase only method starts with higher accuracy but degrades more rapidly as pruning rate increases. Finally, the combined approach exhibits the most robust performance, maintaining superior accuracy until around $0.7$ pruning rate, after which all methods converge to similar performance levels. This suggests that our two-phase design leverages complementary information from both token-level similarity and attention patterns, resulting in more robust and efficient pruning even on challenging cases where the original model fails.
Figure 3: Ablation study comparing different pruning strategies on SPIRAL-H dataset. The plot shows the accuracy of three approaches: first phase only, second phase only, and SpeechPrune. The dotted line shows the accuracy when using the complete, unpruned set of input tokens.
@@ IV-D Generalization Analysis
To evaluate the generalization capability of our method, we test SpeechPrune on both different benchmarks and a different Speech LLM model. For additional benchmarks, we select two representative long-form speech understanding datasets: DREAM-TTS and CN-College-Listen. DREAM-TTS is derived from the text-based dialogue comprehension dataset DREAM [DREAM], converted to speech using state-of-the-art TTS technology with 60 different speakers while maintaining gender consistency as described by [wang2024audiobench]. CN-College-Listen is sourced from WavLLM’s [hu2024wavllm] test set, comprising English listening comprehension questions from China’s national college entrance examinations. For both datasets, we specifically use test samples that exceed 60 seconds in length to evaluate long-form speech understanding capabilities.